OpenStudy (anonymous):

Solve: (x + 1)2 = 9 A. -2, 4 B. 2, -4 C. 2, 4 D. 1, -4

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: if we take the square root of both sides, we get the subsequent equations: \[\large \begin{gathered} x + 1 = 3 \hfill \\ x + 1 = - 3 \hfill \\ \end{gathered} \] since we have: \[\sqrt 9 = \pm 3\]

OpenStudy (michele_laino):

we have to solve those equations

OpenStudy (michele_laino):

for example, let's consider the first equation: x+1=3 If I subtract +1 from both sides, I get: \[\large x + 1 - 1 = 3 - 1\]

OpenStudy (michele_laino):

what is x?

OpenStudy (anonymous):

2?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

that's the first solution. Now let's consider the second equation: x+1=-3 again I subtract +1 from both sides, and I get: \[\large x + 1 - 1 = - 3 - 1\] what is x?

OpenStudy (anonymous):

Oh I just realized that when I pasted the problem it changed the format. (x + 1)^2 = 9 is what it should be. The 2 is an exponent

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Okay?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

SO what next?

OpenStudy (michele_laino):

what is x, from the second equation?

OpenStudy (anonymous):

-4

OpenStudy (michele_laino):

\[\large x + 1 - 1 = - 3 - 1\]

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

2,-4, or B

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Awesome thanks again!

OpenStudy (michele_laino):

thanks!